2x+y=6 and 1/2x+1/4y=1 and 1/3x-1/2=3 and -2x+y=4 I am having so much problems with these kinds of problems. please help.
Answer by stanbon(75887)

2x+y=6 and 1/2x+1/4y=1 and 1/3x-1/2=3 and -2x+y=4
--------
Solve each one for "y" to get it in slope-intercept form:
--------
2x+y=6
Subtract 2x from both sides to get:
y=-2x+6
slope=-2; intercept=6
------------
(1/2x) + (1/4y) =1
Multiply thru by 4xy to get:
2y + x = 4xy
Get the "y" terms together:
4xy-2y=x
Factor out the y to get:
y(4x-2)=x
Solve for y to get:
y=x/(4x-2)
The graph of this is not a line;so, no slope.
---------------
1/3x-1/2=3
Multiply thru by 6x to get:
2 -3x = 18x
21x=2
x=2/21
---------
-2x+y=4
y=2x+4
slope = -2; intercept=4
--------
Cheers,
Stan H.



---------
Cheers,
Stan H.